OpenStudy (anonymous):

Suppose a bag of marbles has 4 green, 2 red, 5 yellow, 1 brown, and 7 blue marbles. What is the probability of picking a green marble, replacing it, and then picking a brown marble?

OpenStudy (anonymous):

@Vocaloid @triciaal

OpenStudy (anonymous):

do you know what the probability of picking a green marble on the first try is?

OpenStudy (anonymous):

@satellite73 4/19

OpenStudy (kash_thesmartguy):

\[\frac{ 4 }{ 19 }*\frac{ 1 }{ 19}=\frac{ 4 }{ 19 }\]